# Adaptive Joint Estimation of Temporal Vertex and Edge Signals

### v2

@@ II-B Signal Processing on graphs and simplicial complexes
Before proceeding to AJVEE, it is necessary to discuss how to reconstruct a static signal $\boldsymbol{x}_{k}$ when only partial observation of the signal is obtainable. With the assumption that $\boldsymbol{x}_{k}$ is a bandlimited signal with frequencies $\mathcal{F}_{k}$, a reconstruction of $\boldsymbol{x}_{k}$ is possible if the conditions $||\mathbf{D}_{k,\bar{\mathcal{S}}}\mathbf{U}_{k,\mathcal{F}}\mathbf{U}_{k,\mathcal{F}}^{T}||_{2}=||\mathbf{U}_{k,\mathcal{F}}\mathbf{U}_{k,\mathcal{F}}^{T}\mathbf{D}_{k,\bar{\mathcal{S}}}||_{2}<1$ and $|\mathcal{F}_{k}|\leq|\mathcal{S}_{k}|$ are satisfied, where $\mathbf{D}_{k,\bar{\mathcal{S}}}=\mathbf{I}-\mathbf{D}_{k,\mathcal{S}}$ [30].
III Adaptive joint vertex-edge estimation
In this section, we propose a joint estimation strategy, targeting the case that the time-varying edge signal influences the time-varying vertex signal, from a simplicial diffusion perspective on a graph. An online simplicial signal estimation procedure named the ALMS-Hodge, based on the estimation error and TSP-based simplicial aggregation, is proposed along with AJVEE to provide adaptive updates in AJVEE.
III-A Vertex regression influenced by time-varying edge signals
$\boldsymbol{x}_{0}[t+1]=\hat{\mathbf{L}}_{0}[t]\boldsymbol{x}_{0}[t]=\hat{\mathbf{B}}_{1}\hat{\mathbf{B}}_{1}^{T}\boldsymbol{x}_{0}[t],$ (4)
where $\hat{\mathbf{L}}_{0}[t]$ is a time-varying regression matrix derived from a graph topology containing a temporal edge signal, and $t$ is the time index. The term $\hat{\mathbf{B}}_{1}$ in equation (4) is
$\hat{\mathbf{B}}_{1}=(\mathbf{B}_{1}\text{diag}(\text{sign}(\boldsymbol{x}_{1}[t])\circ\sqrt{|\boldsymbol{x}_{1}[t]|}))^{T},$ (5)
where $\boldsymbol{x}_{1}[t]$ is the time-varying edge signal, diag$(\boldsymbol{x}_{1}[t])$ forms a diagonal matrix from $\boldsymbol{x}_{1}[t]$ and sign() computes the element-wise sign. We should emphasize that this process is fundamentally different from spatial node aggregation using purely the graph Laplacian $\mathbf{L}_{0}$. The regression matrix $\hat{\mathbf{L}}_{0}[t]$ is formed by mapping time-varying edge signals $\boldsymbol{x}_{1}[t]$ onto the topology and should not simply be perceived as a weighted version of the static graph Laplacian matrix ${\mathbf{L}}_{0}$. By combining $\hat{\mathbf{L}}_{0}[t]$ time-varying edge signals, we result in a time-varying influence of $\boldsymbol{x}_{1}[t]$ on $\boldsymbol{x}_{0}[t]$ in the form of time-varying regression matrix $\hat{\mathbf{L}}_{0}[t]$ that models the change of vertex signal $\boldsymbol{x}_{0}[t]$.
${x}_{i}[t+1]=\sum^{N}_{j=1}\hat{{l}}_{ij}x_{i}[t],$ (6)
where $\hat{{l}}_{ij}$ is the $i^{th}$ row and $j^{th}$ column of $\hat{\mathbf{L}}_{0}[t]$, and $x_{i}[t]$ is the $i^{th}$ element of $\boldsymbol{x}_{0}[t]$. Realizing that in $\hat{\mathbf{L}}_{0}[t]$, $\hat{{l}}_{ij}$ is non-zero only when there is an edge, each summation in equation (6) essentially is calculating the vertex signal at $x_{i}$ minus all the neighboring vertex signals from 1-hop away. If we repeat (6) for all $N_{0}$ vertices, each vertex will be aggregated to its 1-hop adjacent neighbor. If we look at (6) in a diffusion aspect, it essentially means that all the vertices will be propagated to their connected neighbors, meaning that the aggregation done in (6) is equivalent to a diffusion. Notably, the $p^{th}$ power of $\hat{\mathbf{L}}_{0}$ will be a time-varying diffusion in the $p$-hop neighborhood. Following this logic, summing all the terms for $p=1...P$ gives us a combined $1$-hop to $P$-hop regression:
where $\hat{\theta}_{p}$ is the polynomial coefficient. The diffusive nature of this formulation means that missing data on the vertices can be imputed from neighborhood connections as the diffusion progresses. Moreover, when the weights $\hat{\theta}_{p}$ in (7) are defined based on a properly designed simplicial filter, simplicial denoising can be achieved with this formulation; a spectral analysis will be conducted in the next subsection as a guide of how to define the weights. For simplicity, we assume that the topology is known, but the signals on the vertices and edges are noisy or have missing entries at the current time $t$. Preferably, the topology does not change over time. An example of synthetic data based on a real-world traffic network of the above data model is shown in Fig. 1, where time-varying signals are generated on both the graph vertices and the edges.
$t=1$
$t=180$
Figure 1: A graph with time-varying signals on both the vertices and the edges. (left color bar: edges, right color bar: vertices)
Another approach to forming the time-varying vertex regression in (4) is to use spectral filters by setting the SFT in (2) to $k=0$ to form a simplicial (node) filter $\mathbf{\Sigma}_{0}$ to approximate $\hat{\mathbf{L}}_{0}[t]$. Let us represent the filter $\mathbf{\Sigma}_{0}$ using a function $h(\mathbf{\Lambda}_{0}[t])=\mathbf{\Sigma}_{0}$. Then, the regression (4) in terms of the spectral filtering in (3) is
There is a limitation of using spectral methods in practice. The simplicial filter $\mathbf{\Sigma}_{0}$ is defined by the frequency set $\mathcal{F}_{0}$ operating on the frequencies (eigenvalues) of $\hat{\mathbf{L}}_{0}[t]$. This means that spectral filtering operation in (8) requires the eigendecomposition at every time instance when the signals $\boldsymbol{x}_{1}[t]$ changes due to the time-varying formulation in (4). Moreover, the eigendecomposition of the graph Laplacian suffers from high computational cost and numerical instabilities when the graph topology is large [48]. Thus, the spatial formulations are preferred in practice while the spectral formulations provide a signal processing analytical perspective.
Using a shifted Chebyshev polynomial $T_{p}(\mathbf{L}_{0}[t])$ similar to what is shown in [48], we can approximate (8) with the $p$-hop time-varying regression in (7). Assuming that the spatial operations are conducted without considering the time dimension, $h(\hat{\mathbf{L}}_{0}[t])$ in (8) is
$h(\hat{\mathbf{L}}_{0})\approx\sum_{p=0}^{P}\theta_{p}T_{p}(\hat{\mathbf{L}}_{0})=\hat{\mathbf{H}}_{0}[t]=\sum_{p=0}^{P}\hat{\theta}_{p}\hat{\mathbf{L}}_{0}^{p}[t],$ (9)
where $\theta_{p}$ is the coefficients of the Chebyshev polynomial and $P$ is the order of the polynomial. If we explicitly write out the (9) and rearrange the terms, we see that (9) becomes (7).
Even though the formulations in (7) provide diffusion-based vertex regressions under the influence of edge signals, they consider only the diffusion perspective but not the online estimation perspective. In the next sections, we will improve this formulation by considering the temporal changes to achieve adaptive estimation.
III-B The ALMS-Hodge procedure
@@
Before completing AJVEE, we would like to propose a novel adaptive filtering algorithm named the ALMS-Hodge algorithm for online time-varying simplicial signal prediction on the $k$-simplices. Let us consider a graph $\mathcal{G}$ as a simplicial complex $\mathcal{K}$ containing time-varying signal $\boldsymbol{x}_{k}[t]$. In classical signal processing, a simple data model can be used to represent the time-varying dynamics of signals:
@@
$\mathbf{J}(\hat{\boldsymbol{x}}_{k}[t])=\mathbb{E}||\boldsymbol{y}_{k}[t]-{\mathbf{D}_{k,\mathcal{S}}}\mathbf{U}_{k,\mathcal{F}}\mathbf{U}_{k,\mathcal{F}}^{T}\hat{\boldsymbol{x}}_{k}[t]||^{2}_{2}+g(\hat{\boldsymbol{x}}_{k}[t]),$ (12)
where $\hat{\boldsymbol{x}}_{k}[t]$ is the estimated bandlimited signal with the same frequency bands as $\boldsymbol{x}_{k}[t]$, and $g(\hat{\boldsymbol{x}}_{k}[t])$ is a task-specific simplicial aggregation. Combining bandlimitedness of $\hat{\boldsymbol{x}}_{k}[t]$ and the idempotent property of $\mathbf{D}_{k,\mathcal{S}}$, the term $\boldsymbol{y}_{k}[t]-{\mathbf{D}_{k,\mathcal{S}}}\mathbf{U}_{k,\mathcal{F}}\mathbf{U}_{k,\mathcal{F}}^{T}\hat{\boldsymbol{x}}_{k}[t]$ in (12) is written as $\mathbf{D}_{k,\mathcal{S}}(\boldsymbol{y}_{k}[t]-\mathbf{U}_{k,\mathcal{F}}\mathbf{U}_{k,\mathcal{F}}^{T}\mathbf{D}_{k,\mathcal{S}}\hat{\boldsymbol{x}}_{k}[t])$. We formulate (12) as a convex least mean squares optimization problem
@@
In AJVEE, an essential step in online estimation is to effectively represent change $\Delta_{k}[t]$ on each simplex using TSP. Looking at (13), we can represent $\Delta_{k}[t]$ as the inverse direction of causes the largest error to minimize the error, which is obtained by the (negative) gradient of $J(\hat{\boldsymbol{x}}_{k}[t])$. By letting $\Delta_{k}[t]=-\frac{\partial\mathbf{J}(\hat{\boldsymbol{x}}_{k}[t])}{\partial\hat{\boldsymbol{x}}_{k}[t]}$ in (10), an adaptive update function based on (10) for online simplicial signal estimation on order $k$ using the Hodge Laplacian can be derived as
$\begin{split}\hat{\boldsymbol{x}}_{k}[t+1]&=\hat{\boldsymbol{x}}_{k}[t]+\Delta_{k}[t]=\hat{\boldsymbol{x}}_{k}[t]-\frac{\partial\mathbf{J}(\hat{\boldsymbol{x}}_{k}[t])}{\partial\hat{\boldsymbol{x}}_{k}[t]}\\ &=\hat{\boldsymbol{x}}_{k}[t]+2h({\mathbf{L}}_{k})\boldsymbol{e}_{k}[t]-\mathbf{R}_{k}\hat{\boldsymbol{x}}_{k}[t],\end{split}$ (14)
where $\boldsymbol{e}_{k}[t]=\mathbf{D}_{k,\mathcal{S}}(\boldsymbol{y}_{k}[t]-\hat{\boldsymbol{x}}_{k}[t])$ is the estimation error of all sampled part of the signal, and $\mathbf{R}_{k}\hat{\boldsymbol{x}}_{k}[t]=\frac{\partial g(\hat{\boldsymbol{x}}_{k}[t])}{\partial\hat{\boldsymbol{x}}_{k}[t]}$. Here, for $h(\hat{\mathbf{L}}_{k}$), we will assume a general formulation and use again the Chebyshev approximation to obtain a spatial update:
$h({\mathbf{L}}_{k})=\mathbf{U}_{k}\mathbf{\Sigma}_{k}\mathbf{U}_{k}^{T}\approx\mathbf{H}_{k}=\sum_{p=0}^{P}\hat{\theta}_{p}{\mathbf{L}}_{k}^{p},$ (15)
where the design of the filter $\mathbf{\Sigma}_{k}$ controls reconstruction properties. Step size parameters $\mu_{k}$ and $r_{k}$ are assigned to control the magnitude of the updates as conventionally seen in classical adaptive filters [50]. The update function of our ALMS-Hodge is derived as
@@
where $\mu_{k}$ is the step size parameter and $r_{k}$ is the weight parameter for the aggregation. The frequencies are defined as larger eigenvalues corresponding to high frequencies, so we can define $\mathbf{\Sigma}_{k}$ as low-pass, high-pass, or band-limited filters similar to how these types of filters are designed in classical signal processing. Viewed in the spectral domain, each update term $\mu_{k}\mathbf{H}_{k}\boldsymbol{e}_{k}[t]$ is applying the filter $\mathbf{\Sigma}_{k}$ to the SFT of the error component $\boldsymbol{e}_{k}[t]$. As a result, at each time instance $t$, the term $\mathbf{H}_{k}\boldsymbol{e}_{k}[t]$ in (16) ensures the update of the ALMS-Hodge is time-varying and is adaptively generated in the direction opposite to the error. It can be shown that the ALMS-Hodge in (16) converges, for the details please check Appendix A.
If the term $r_{k}\mathbf{R}_{k}\hat{\boldsymbol{x}}_{k}[t]$ is ignored as seen in previous adaptive GSP algorithms, the ALMS-Hodge will update solely based on the errors $\boldsymbol{e}_{k}[t]$ but not utilizing the signals $\boldsymbol{x}_{k}[t]$ itself. Thus, it is advantageous if we define aggregation terms based on the underlying signal to enforce additional restrictions or enhancements. For example, if we want to emphasize the property $\mathbf{B}_{k}^{T}\mathbf{B}_{k}\hat{\boldsymbol{x}}_{k}[t]=0$ on the lower adjacency of $\mathcal{X}_{k}$, then we can define an aggregation term $g(\hat{\boldsymbol{x}}_{k}[t])=||\mathbf{B}_{k}\hat{\boldsymbol{x}}_{k}[t]||^{2}_{2}$ in the cost function (12). Following the optimization steps in (14), the resulting aggregation $\mathbf{R}_{k,l}\hat{\boldsymbol{x}}_{k}[t]$ in the update equation (23) is
@@
where $r_{k}=2\nu_{k}$ is the weight for the aggregation. Similarly, we can define the upper adjacent aggregation $\mathbf{R}_{k,u}\hat{\boldsymbol{x}}_{k}[t]$ using the same approach by setting $h(\hat{\boldsymbol{x}}_{k}[t])=||\mathbf{B}_{k+1}^{T}\hat{\boldsymbol{x}}_{k}[t]||^{2}_{2}$:
@@
The terms $\mathbf{B}_{k}^{T}\boldsymbol{x}_{k-1}[t]$ and $\mathbf{B}_{k+1}\boldsymbol{x}_{k+1}[t]$ are the signal components that can be induced from $\mathcal{X}_{k-1}$ and $\mathcal{X}_{k+1}$ signals respectively. The term $\boldsymbol{x}_{k,H}[t]$ in (19) is the harmonic component corresponding to the portion of the signal that cannot be induced from $\boldsymbol{x}_{k-1}$ or $\boldsymbol{x}_{k+1}$. It is worth mentioning that these three components are orthogonal [30]. Inspired by the Hodge decomposition, if we revisit the regression/diffusion defined in (7), at each time instance $t$, we can also define aggregation based on the boundary and coboundary. For example, if we want to aggregate the induced components from $\mathcal{X}_{k-1}$ onto $\mathcal{X}_{k}$, we can use the boundary aggregation
@@
Note that the above four aggregations can be linearly combined into a single term $\mathbf{R}_{k}\hat{\boldsymbol{x}}_{k}[t]$ in AJVEE to achieve desired effects.
@@
III-C Joint estimation of vertex and edge signals

### v3

@@ II-B Signal Processing on graphs and simplicial complexes
Before proceeding to AJVEE, it is necessary to discuss how to reconstruct a static signal $\boldsymbol{x}_{k}$ when only partial observation of the signal is obtainable. With the assumption that $\boldsymbol{x}_{k}$ is a bandlimited signal with frequencies $\mathcal{F}_{k}$, a reconstruction of $\boldsymbol{x}_{k}$ is possible if the conditions $\|\mathbf{D}_{k,\bar{\mathcal{S}}}\mathbf{U}_{k,\mathcal{F}}\mathbf{U}_{k,\mathcal{F}}^{T}\|_{2}=\|\mathbf{U}_{k,\mathcal{F}}\mathbf{U}_{k,\mathcal{F}}^{T}\mathbf{D}_{k,\bar{\mathcal{S}}}\|_{2}<1$ and $|\mathcal{F}_{k}|\leq|\mathcal{S}_{k}|$ are satisfied, where $\mathbf{D}_{k,\bar{\mathcal{S}}}=\mathbf{I}-\mathbf{D}_{k,\mathcal{S}}$ [16].
III The ALMS-Hodge
@@
Before discussing the AJVEE framework, we would like to introduce a novel adaptive filtering algorithm named the ALMS-Hodge for online time-varying simplicial signal prediction on the $k$-simplices. Let us consider a graph $\mathcal{G}$ as a simplicial complex $\mathcal{K}$ containing time-varying signal $\boldsymbol{x}_{k}[t]$. In classical signal processing, a simple data model can be used to represent the time-varying dynamics of signals:
@@
$\mathbf{J}(\hat{\boldsymbol{x}}_{k}[t])=f(\hat{\boldsymbol{x}}_{k}[t])+g(\hat{\boldsymbol{x}}_{k}[t]),$ (11)
where $f(\hat{\boldsymbol{x}}_{k}[t])=\mathbb{E}\|\boldsymbol{y}_{k}[t]-{\mathbf{D}_{k,\mathcal{S}}}\mathbf{U}_{k,\mathcal{F}}\mathbf{U}_{k,\mathcal{F}}^{T}\hat{\boldsymbol{x}}_{k}[t]\|^{2}_{2}$ is the $l_{2}$ cost that we are aiming to optimize, $\hat{\boldsymbol{x}}_{k}[t]$ is the estimation of the ground truth signal ${\boldsymbol{x}}_{k}[t]$, and $g(\hat{\boldsymbol{x}}_{k}[t])$ is a task-specific regularization in terms of simplicial aggregation that we will be deriving later. At $t=0$, the estimation $\boldsymbol{x}_{k}[0]$ is the initialization of the ALMS-Hodge. Here we assume that both the estimated signal $\hat{\boldsymbol{x}}_{k}[t]$ and $\boldsymbol{x}_{k}[t]$ are bandlimited and share the same frequency component. This consistency is ensured because both $\boldsymbol{x}_{k}[t]$ and $\hat{\boldsymbol{x}}_{k}[t]$ use the same topology, meaning that the SFT and frequency components are defined by the same Hodge Laplacian. The filter design of $\mathbf{\Sigma}_{k,\mathcal{F}}$ can follow a universal filter such as the smoothness-based low-pass filter [25]. Another option we can choose is to use the SFT to acquire the spectrum directly from the historical data or estimate the spectrum from the noisy observation. We formulate (11) as a convex least mean squares optimization problem similar to GSP on the vertices [4]:
@@
In ALMS-Hodge, an essential step in online estimation is to effectively represent the change $\Delta_{k}[t]$ on each simplex using TSP techniques. Looking at (12), we can represent $\Delta_{k}[t]$ as the change between two-time points, which is obtained by the (negative) gradient of $J(\hat{\boldsymbol{x}}_{k}[t])$. By letting $\Delta_{k}[t]=-\frac{\partial\mathbf{J}(\hat{\boldsymbol{x}}_{k}[t])}{\partial\hat{\boldsymbol{x}}_{k}[t]}$ in (9), an adaptive update function based on (9) for online simplicial signal estimation on order $k$ using the Hodge Laplacian can be derived as
$\hat{\boldsymbol{x}}_{k}[t+1]=\hat{\boldsymbol{x}}_{k}[t]+\Delta_{k}[t]=\hat{\boldsymbol{x}}_{k}[t]-\frac{\partial\mathbf{J}(\hat{\boldsymbol{x}}_{k}[t])}{\partial\hat{\boldsymbol{x}}_{k}[t]}.$ (13)
Combining bandlimitedness of $\hat{\boldsymbol{x}}_{k}[t]$ and the idempotent property of $\mathbf{D}_{k,\mathcal{S}}$, the term $\boldsymbol{y}_{k}[t]-{\mathbf{D}_{k,\mathcal{S}}}\mathbf{U}_{k,\mathcal{F}}\mathbf{U}_{k,\mathcal{F}}^{T}\hat{\boldsymbol{x}}_{k}[t]$ can be written as $\mathbf{D}_{k,\mathcal{S}}(\boldsymbol{y}_{k}[t]-\mathbf{U}_{k,\mathcal{F}}\mathbf{U}_{k,\mathcal{F}}^{T}\mathbf{D}_{k,\mathcal{S}}\hat{\boldsymbol{x}}_{k}[t])$. The update function of our ALMS-Hodge is derived as
@@
where $-2\mathbf{H}_{k}{\boldsymbol{e}}_{k}[t]=\frac{\partial f(\hat{\boldsymbol{x}}_{k}[t])}{\partial\hat{\boldsymbol{x}}_{k}[t]}$ and $\mathbf{R}_{k}\hat{\boldsymbol{x}}_{k}[t]=\frac{\partial g(\hat{\boldsymbol{x}}_{k}[t])}{\partial\hat{\boldsymbol{x}}_{k}[t]}$. The term $\boldsymbol{e}_{k}[t]=\mathbf{D}_{k,\mathcal{S}}(\boldsymbol{y}_{k}[t]-\hat{\boldsymbol{x}}_{k}[t])$ in (14) is the estimation error of all sampled part of the signals. Here, we can convert the spectral update into a general formulation of spatial update and use again the Chebyshev approximation in (7) to obtain a spatial update:
$\mathbf{H}_{k}=\mathbf{U}_{k,\mathcal{F}}\mathbf{U}_{k,\mathcal{F}}^{T}\approx{\theta_{0}}+\sum_{p=1}^{P}\theta_{p}T_{p}(\mathbf{L}_{k}).$ (15)
Step size parameters $\mu_{k}$ and $r_{k}$ are assigned to control the magnitude of the update terms as conventionally seen in classical adaptive filters [40]. The frequencies are defined as larger eigenvalues corresponding to high frequencies, so we can define $\mathbf{\Sigma}_{k}$ as low-pass, high-pass, or band-limited filters similar to how these types of filters are designed in classical signal processing. Viewed in the spectral domain, each update term $\mu_{k}\mathbf{H}_{k}\boldsymbol{e}_{k}[t]$ is applying the filter $\mathbf{\Sigma}_{k}$ to the SFT of the error component $\boldsymbol{e}_{k}[t]$. As a result, at each time instance $t$, the term $\mathbf{H}_{k}\boldsymbol{e}_{k}[t]$ in (14) ensures the update of the ALMS-Hodge is time-varying and is adaptively generated in the direction opposite to the error. It can be shown that the ALMS-Hodge in (14) converges, for the details please check Appendix A.
If the term $r_{k}\mathbf{R}_{k}\hat{\boldsymbol{x}}_{k}[t]$ is ignored as seen in previous adaptive GSP algorithms, the ALMS-Hodge will update solely based on the errors $\boldsymbol{e}_{k}[t]$ but not utilizing the signals $\boldsymbol{x}_{k}[t]$ itself. Thus, it is advantageous if we define aggregation terms based on the underlying signal to enforce additional restrictions or enhancements. For example, if we want to emphasize the property $\mathbf{B}_{k}^{T}\mathbf{B}_{k}\hat{\boldsymbol{x}}_{k}[t]=0$ on the lower adjacency of $\mathcal{X}_{k}$, then we can define an aggregation term $g(\hat{\boldsymbol{x}}_{k}[t])=\|\mathbf{B}_{k}\hat{\boldsymbol{x}}_{k}[t]\|^{2}_{2}$ in the cost function (11). Following the optimization steps in (13), the resulting aggregation $\mathbf{R}_{k,l}\hat{\boldsymbol{x}}_{k}[t]$ in the update equation (14) is
@@
where $r_{k}=2\nu_{k}$ is the weight for the aggregation. Similarly, we can define the upper adjacent aggregation $\mathbf{R}_{k,u}\hat{\boldsymbol{x}}_{k}[t]$ using the same approach by setting $h(\hat{\boldsymbol{x}}_{k}[t])=\|\mathbf{B}_{k+1}^{T}\hat{\boldsymbol{x}}_{k}[t]\|^{2}_{2}$:
@@
The terms $\mathbf{B}_{k}^{T}\boldsymbol{x}_{k-1}[t]$ and $\mathbf{B}_{k+1}\boldsymbol{x}_{k+1}[t]$ are the signal components that can be induced from $\mathcal{X}_{k-1}$ and $\mathcal{X}_{k+1}$ signals respectively. The term $\boldsymbol{x}_{k,H}[t]$ in (18) is the harmonic component corresponding to the portion of the signal that cannot be induced from $\boldsymbol{x}_{k-1}$ or $\boldsymbol{x}_{k+1}$. It is worth mentioning that these three components are orthogonal [16]. Remember in Section II, the upper adjacency, lower adjacency, boundary, and coboundary of $\mathcal{X}_{k}$ can be defined using $\mathbf{B}_{k}$ and $\mathbf{B}_{k+1}$. Inspired by the Hodge decomposition, at each time instance $t$, we can define regularization terms in terms of simplicial aggregation based on the boundary and coboundary. For example, if we want to aggregate the induced components from $\mathcal{X}_{k-1}$ onto $\mathcal{X}_{k}$, we can use the boundary aggregation
@@
Note that the above four aggregations can be linearly combined in AJVEE to achieve the desired effects.
@@
There are three main advantages of the ALMS-Hodge. First, unlike the classical adaptive filter that learns the filters while making online estimation, our proposed ALMS-Hodge relies on assumptions such as spatial smoothness and bandlimitedness of the simplicial signal, then predefine universal TSP filters such as the low-pass filter. This means that the ALMS-Hodge will not be restricted to a specific order of simplicial signals. Additionally, not needing to learn the filters from the data reduces the complexity of the algorithm and simplifies the implementation in practice. Lastly, the ALMS-Hodge directly uses topological information of the data, which allows the missing simplicial signals to be estimated along with the temporal prediction. This is a unique characteristic of GSP and TSP algorithms that is not seen in classical signal processing.
$t=1$
$t=180$
Figure 1: The Sioux Falls network with time-varying signals on both the vertices and the edges. (left color bar: edges, right color bar: vertices)
IV Adaptive joint vertex-edge estimation
In this section, the AJVEE is proposed to target the case that the time-varying edge signal influences the time-varying vertex signal, from a simplicial diffusion perspective on a graph. The ALMS-Hodge will be used to facilitate the operation of AJVEE.
IV-A Vertex regression influenced by time-varying edge signals
$\boldsymbol{x}_{0}[t+1]=\hat{\mathbf{L}}_{0}[t]\boldsymbol{x}_{0}[t]={\mathbf{B}}_{1}\text{diag}(|\boldsymbol{x}_{1}[t]|){\mathbf{B}}_{1}^{T}\boldsymbol{x}_{0}[t],$ (22)
where $\hat{\mathbf{L}}_{0}[t]$ is a time-varying regression matrix derived from a graph topology containing a temporal edge signal, and $t$ is the time index. We should emphasize that this process is fundamentally different from spatial node aggregation using purely the graph Laplacian $\mathbf{L}_{0}$. The regression matrix $\hat{\mathbf{L}}_{0}[t]$ is formed by mapping time-varying edge signals $\boldsymbol{x}_{1}[t]$ onto the topology and should not simply be perceived as a static weighted version of the graph Laplacian matrix ${\mathbf{L}}_{0}$. By combining $\hat{\mathbf{L}}_{0}[t]$ time-varying edge signals, we result in a time-varying influence of $\boldsymbol{x}_{1}[t]$ on $\boldsymbol{x}_{0}[t]$ in the form of time-varying regression matrix $\hat{\mathbf{L}}_{0}[t]$ that models the change of vertex signal $\boldsymbol{x}_{0}[t]$.
${x}_{i}[t+1]=\sum^{N}_{j=1}\hat{{l}}_{ij}x_{i}[t],$ (23)
where $\hat{{l}}_{ij}$ is the $i^{th}$ row and $j^{th}$ column of $\hat{\mathbf{L}}_{0}[t]$, and $x_{i}[t]$ is the $i^{th}$ element of $\boldsymbol{x}_{0}[t]$. Realizing that in $\hat{\mathbf{L}}_{0}[t]$, $\hat{{l}}_{ij}$ is non-zero only when there is an edge, each summation in equation (23) essentially is calculating the vertex signal at $x_{i}$ minus all the neighboring vertex signals from 1-hop away. When repeating (23) for all $N_{0}$ vertices, each vertex will be aggregated to its 1-hop adjacent neighbor. If we look at (23) in a diffusion aspect, it essentially means that all the vertices will be propagated to their connected neighbors, meaning that the aggregation done in (23) is equivalent to a diffusion.
Following the above logic, the $p^{th}$ power of $\hat{\mathbf{L}}_{0}$ will be a time-varying diffusion in the $p$-hop neighborhood. Summing all the terms for $p=0...P$ gives us a combined $1$-hop to $P$-hop regression:
where $\hat{\theta}_{p}$ is introduced as polynomial coefficients. The diffusive nature of this formulation means that missing data on the vertices can be imputed from neighborhood connections as the diffusion progresses. Moreover, when the weights $\hat{\theta}_{p}$ in (24) are defined based on a properly designed simplicial filter, simplicial denoising can be achieved with this formulation; a spectral analysis will be conducted in the next subsection as a guide of how to define the weights. The expression in (24) can be realized by the Chebyshev approximation shown in (7). If we plug $\hat{\mathbf{L}}_{0}[t]$ into (7) and rearrange the terms, the resulting expression becomes a spatial diffusion represented by $\hat{\mathbf{L}}_{0}[t]$:
${\theta_{0}}+\sum_{p=1}^{P}\theta_{p}T_{p}(\hat{\mathbf{L}}_{0}[t])=\sum_{p=0}^{P}\hat{\theta}_{p}\hat{\mathbf{L}}_{0}^{p}[t]=\hat{\mathbf{H}}_{0}[t],$ (25)
where now $\hat{\theta}_{p}$ is the polynomial coefficient instead of $\theta_{p}$.
Another approach to forming the time-varying vertex regression in (22) is to use spectral filters by setting the SFT in (2) to $k=0$ to form a simplicial (node) filter $\mathbf{\Sigma}_{0,\mathcal{F}}$ to approximate $\hat{\mathbf{L}}_{0}[t]$. Let us represent the filter $\mathbf{\Sigma}_{0,\mathcal{F}}$ using a function $h(\mathbf{\Lambda}_{0}[t])=\mathbf{\Sigma}_{0,\mathcal{F}}$. Then, the regression (22) in terms of the spectral filtering in (3) is
Applying the logic of using a shifted Chebyshev polynomial $T_{p}(\mathbf{L}_{0}[t])$ similar to what is shown in (6), we can approximate (26) with the $p$-hop time-varying regression in (24):
$h(\hat{\mathbf{L}}_{0})\approx{\theta_{0}}+\sum_{p=1}^{P}\theta_{p}T_{p}(\hat{\mathbf{L}}_{0}[t])=\sum_{p=0}^{P}\hat{\theta}_{p}\hat{\mathbf{L}}_{0}^{p}[t]=\hat{\mathbf{H}}_{0}[t].$ (27)
Equation (27) establishes the relationship between the spectral domain and the spatial domain, illustrating the underlying filter $\mathbf{\Sigma}_{0,\mathcal{F}}$ can be applied from two different perspectives.
There is a limitation of using spectral methods in practice. The simplicial filter $\mathbf{\Sigma}_{0,\mathcal{F}}$ is defined by the frequency set $\mathcal{F}_{0}$ operating on the frequencies (eigenvalues) of $\hat{\mathbf{L}}_{0}[t]$. This means that spectral filtering operation in (26) requires the eigendecomposition at every time instance when the signals $\boldsymbol{x}_{1}[t]$ changes due to the time-varying formulation in (22). Moreover, the eigendecomposition of the graph Laplacian suffers from high computational cost and numerical instabilities when the graph topology is large [43]. The same limitation can be found in TSP for the Hodge-Laplacian [24]. Thus, the spatial formulations are preferred in practice while the spectral formulations provide a signal processing analytical perspective.
For simplicity, we assume that the topology is known, but the signals on the vertices and edges are noisy or have missing entries at the current time $t$. Preferably, the topology does not change over time. An example of synthetic data based on a real-world traffic network of the above data model is shown in Fig. 1, where time-varying signals are generated on both the graph vertices and the edges. Even though the formulations in (24) provide diffusion-based vertex regressions under the influence of edge signals, they consider only the diffusion perspective but not the online estimation perspective. We will improve this formulation by considering the temporal changes to achieve adaptive estimation.
IV-B Joint estimation of vertex and edge signals